For each of the following, fill in the table of values.
a. f(X) = lxl
(It has all of the X values listed. They are as follows: -3, -2, -1, zero, one, two, three. I have to fill in the, that is the output: f(X). How do I do this?
b. F(X) = x^3
C. f(X) = 2^x

Answer by Alan3354(69443)

F(x) = |x|
-----
Whevever you see an x, sub the value.
F(-3) = |-3| = 3
F(-2) = |-2| = 2
Etc.
-------
F(x) = x^3
F(-3) = (-3)^3 = -27
.
.
.
F(0) = 0
F(1) 1^3 = 1
Etc
============
F(x) = 2^x
F(-3) = 2^(-3) = 1/8
.
.
F(0) = 2^0 = 1
F(3) = 2^3 = 8
----------------------------
People get nervous about functions.
It's just substitution.